# Incentivize Contribution and Learn Parameters Too: Federated Learning with Strategic Data Owners

### v3

@@ 3 Learning with arbitrary effective costs
Using the update rules of Algorithm˜1 given by Equation˜3 and Equation˜4, we get
@@
1
Output: $w^{T}$
2
3for $t=0$ to T1T-1 do
4 Center broadcasts $w^{t},s^{t}$
5
6 foreach agent $i\in N$ in parallel do
7    $s^{t+1}_{i}=s^{t}_{i}+\gamma[g(w^{t},s^{t},\mu^{t})]_{i}$
9      Send $s_{i}^{t+1},d_{i}^{t+1}$ to the center
10
11
13
14
15return $w^{T}$
16
Algorithm 1 Updated Parameter Best Response Dynamics (UPBReD)
The theorem provides a guarantee of convergence of Algorithm˜1 for any arbitrary initial condition $(w^{0},s^{0})$. It needs to run for a minimum number of iterations $T_{0}(w^{0},s^{0})$, with appropriate parameters of $\gamma$ and $\eta$ that govern the gradient ascent rates of the agents’ and center’s objectives respectively. Note that for a better convergence, i.e., a smaller $\epsilon$, the algorithm needs to run longer. The parameter $W$ determines the contraction factor of the recurrence of $\|g(w^{t},s^{t})\|_{2}+\|\tilde{g}(w^{t},s^{t})\|_{2}$ and is chosen to be smaller than unity by choosing $\gamma$ and $\eta$ appropriately. The following example shows an example of such a feasible region.
Example 1 (Feasible region of $\gamma$ for Theorem˜1 to hold).
(a) $\tilde{P}=1$
(b) $\lambda=1.1$
Remarks. The step sizes $\gamma$ and $\nu$ in our analysis scale inversely with $m$ and $n$, and certain restrictions on $\lambda$ and $\tilde{\lambda}$ may appear restrictive. In Section 4, we address these issues and propose a learning algorithm with step sizes independent of $m$ and $n$.
While agents strategically choose $s_{i}$ to maximize their utility, we assume truthful reporting of these choices. In UPBReD, agents report $s_{i}^{t}$ and $d_{i}^{t}$ in each round. The next subsection introduces payments to ensure truthful reporting of $s_{i}^{t}$ and $d_{i}^{t}$.
3.1 Truthful elicitation with payment rules
So far, we have assumed that the agents report their type truthfully in every round, i.e. if an agent is training on $s_{i}^{t}$ amount of data at time $t$ the agent reports $\theta_{i}=(s_{i}^{t},d_{i}^{t})$ truthfully. However, an agent can misreport their strategy and corresponding gradient update to improve their utility as observed in the example below.
Example 2.
We perform some experiments on the CIFAR data set to show that agents have an incentive to misreport. Consider $2$ agents each having a set of $480$ data points. $u_{i}=56-L(w,s)-c_{i}(s_{i})$, where $L(w,s)$ is the cross entropy loss computed on the dataset. The costs for the agents are $c_{1}(s_{1})=0.2s_{i},c_{2}(s_{2})=0.1s_{2}$. When agents truthfully report $s_{i},d_{i}$ in UPBReD they observe a social welfare of $15.5$, with $s_{1}*=270,s_{2}*=165$ this yields a utility $u_{1}=510$. When agent misreports $s_{1}^{t}=240$ and reports $d_{1}^{t}=0_{m}$ in every training round yielding a utility of $u_{i}=64$. Clearly agent $1$ is better off by misreporting their type. ∎
A well established goal in social choice theory is to choose the allocation that maximizes the social welfare (Green and Laffont, 1977). In our setting this maximizing allocation for a given type profile $\theta^{t}$ at round $t$ will be
$a_{*}({\theta}^{t})=\operatorname*{\arg\!\max}_{w\in\mathbb{R}^{m}}{\sum_{i\in N}v_{i}(w,s^{t}})$ (12)
However, in our federated learning setting, we cannot compute this welfare maximizing allocation. Instead, we use $a_{w}^{t}$, which iteratively takes gradient ascent steps towards the welfare maximizing allocation as shown in Algorithm˜1. The initial allocation is chosen as $w^{0}$. And future allocations are computed iteratively as $a_{w}^{t}(\theta^{t})=w^{t}(s^{t},d^{t})=w^{t-1}+\frac{\eta}{n}\sum_{i\in N}d_{i}^{t}$. Due to this difference, we need an additional assumption for truthfulness. We assume the difference between the welfare at our computed allocation and that of the maximum at truthful reports is bounded by a constant.
Assumption 2.
Consider the following payment rule:
where $\phi$ is given by,
$\textstyle\phi\left(\hat{v}^{t}_{i},v_{i}(a_{w}^{t}(\hat{\theta}^{t}),\hat{s}^{t})\right)$ $\textstyle=0\quad\text{if }\hat{v}^{t}_{i}=v_{i}(a_{w}^{t}(\hat{\theta}^{t}),\hat{s}^{t})$
$\textstyle=\Gamma+(\hat{v}^{t}_{i}-v_{i}(a_{w}^{t}(\hat{\theta}^{t}),\hat{s}^{t}))^{2}\quad\text{otherwise}$
  1. 1.
  2. 2.
  3. 3.
    Agents share their type as $\hat{\theta}_{i}^{t}=(\hat{s}_{i}^{t},\hat{d}_{i}^{t})$ which can be different from their true type. The center computes an allocation using the reported types using $a_{w}^{t}$ consistent with Equation˜4.
  4. 4.
[…]
  5. 5.
With the given allocation rule $a_{w}^{t}$ and payment rule $p$ described in Equation˜13 for a true type and valuation profile $\theta^{t},v^{t}$, a reported type and valuation profile $\hat{\theta}^{t},\hat{v}^{t}$, the utility realised by agent $i$ at round $t$ is given by $u_{i}(\hat{\theta}^{t},\hat{v}^{t}|\theta^{t},v^{t})=v_{i}(a_{w}^{t}(\hat{\theta}^{t}),s^{t})-z_{i}(s_{i}^{t},\hat{s}_{-i}^{t})+p_{i}^{*}(\hat{\theta}^{t},\hat{v}_{i}^{t})$. The algorithm is detailed in Algorithm˜2.
@@ 17     Send $\hat{v}_{i}^{t}=\hat{a}_{i}$
19 Center computes payments using Equation˜13 at $w^{t}$
@@
With this algorithm, we can show the following result.
Theorem 2.
Under ˜2, Algorithm˜2 is EPIC in every round.
Proof.
@@
Expanding the terms as defined in Equation˜13
@@
As social welfare is maximized by Equation˜12
@@
3.2 Application of Theorem˜1: effective costs are only personal
Example 3 (NE different from socially optimal).
So, our objective in this paper is to allow monetary transfer among the agents so that we get the best of both worlds: (a) We achieve convergence to a Nash equilibrium of $(w^{\texttt{{OPT}}},s^{\max})$, where all agents to contribute their entire data and center learns the optimal parameter, and (b) the transfers are budget balanced, i.e., the center does not accumulate any money – it is used just to realign the agents’ utilities to reach the desired Nash equilibrium. We discuss this mechanism in the following section.
4 Welfare maximization at Nash equilibrium
In this section, we consider the scenario where the center can make a payment $p_{i}(s_{i},s_{-i})$ to agent $i$ to alter its utility function. Therefore, the effective cost of agent $i$ becomes $z_{i}(s_{i},s_{-i})=c_{i}(s_{i})-p_{i}(s_{i},s_{-i})$. Note that, the payment can be either positive or negative, which determines whether the agent is subsidized or taxed respectively. It is reasonable to expect that the data contributors of the federated learning process are subsidized for their data contribution while the data consumers are charged payment for obtaining the learned parameter. We assume that the derivatives of the valuation $v_{i}$ and cost function $c_{i}$ are bounded. We choose the following payment function:
$p_{i}(s)=\beta\left(s_{i}-\frac{\textstyle 1}{\textstyle n-1}\sum_{j\neq i}s_{j}\right),$ (14)
where $\beta$ is a parameter of choice. Note that this payment mechanism is budget balanced by design, since $\sum_{i\in N}p_{i}(s_{i},s_{-i})=0,\forall s_{i}\in S_{i},s_{-i}\in S_{-i}$. With payment, the utility of agent $i$ becomes:
$u_{i}(w,s)=v_{i}(w,s)-c_{i}(s_{i})+\beta(s_{i}-\frac{\textstyle 1}{\textstyle n-1}\sum_{j\neq i}s_{j}).$ (15)
In this section, we are also interested in the quality of the NE in terms of social welfare and want to reach the desired NE where $s_{i}^{*}=s_{i}^{\max},\forall i\in N$ and $w^{*}=w^{\texttt{{OPT}}}$.
4.1 Convergence to the NE $(w^{\textnormal{OPT}},s^{\max})$
Assumption 3.
Note that the above assumptions are pretty mild. Previous works assume even stronger assumptions (like convexity) on $v_{i}$ and $c_{i}$ (see (Murhekar et al., 2023)). We show that the above bounds on the derivatives of $v_{i}$ and $c_{i}$ are sufficient to claim our result.
Lemma 1 (Increasing utility).
Suppose Assumption 3 holds and we use the payment function given by Equation˜14. Then, the utility of every agent is increasing in its own data contribution, i.e., $\frac{\partial}{\partial s_{i}}u_{i}(w,s_{i},s_{-i})>0,\forall s_{i}\in(0,s_{i}^{\max}),\forall i\in N,\text{ and }\forall w$ provided $\beta>\zeta+\tau$.
@@ 4.1 Convergence to the NE $(w^{\textnormal{OPT}},s^{\max})$
From Equation˜15, we get the derivative of the utility function for $s_{i}\in(0,s_{i}^{\max})$ (note that this holds only for the interior of $S_{i}$, the derivative at the boundaries are zero by definition of function $g$ in Section˜3) to be $g(w,s,\mu)]_{i}=\frac{\partial}{\partial s_{i}}u_{i}(w,s)=\frac{\partial}{\partial s_{i}}v_{i}(w,s)-c_{i}^{\prime}(s_{i})+\beta$, where $c_{i}^{\prime}(s_{i})=\frac{d}{ds_{i}}c_{i}(s_{i})$. Using Assumption 3 and choosing in Equation˜14 the parameter $\beta>\zeta+\tau$, we have
@@
Definition 4 ($M$-smoothness).
Definition 5 ($\nu$-strictly convex).
1
2
Output: $w^{T}$
Center broadcasts $w^{0},s^{0}$, set $t=0$
/* Begin phase 1 */
3 while $s\neq s^{\max}$ do
4 Center broadcasts $s^{t}$
5 for agent $i\in N$ in parallel do
6    $s_{i}^{t+1}=s_{i}^{t}+\gamma[g(w^{0},s^{t},\mu^{t})]_{i}$
7      Send $s_{i}^{t+1}$ to the center
8
 $t=t+1$
 /* End phase 1 */
9
/* Begin phase 2 */
10 for $t=0$ to $T$ do
11 Center broadcasts $w^{t}$
12 for agent $i\in N$ in parallel do
14      Send $d_{i}^{t+1}$ to the center
15
return $w^{T}$
/* End phase 2 */
17
Algorithm 3 Two Phase Updated Parameter Best Response Dynamics (2P-UPBReD)
We propose a two-phase algorithm given by Algorithm˜3. The algorithm, in the first phase, incentivizes the agents to contribute $s^{\max}$, and in the second phase, converges to $w^{\texttt{{OPT}}}$. Note that in the first phase of Algorithm˜3, every agent runs a gradient ascent step. Thanks to the increasing utility property (Lemma 1), we have an increasing sequence of $s_{i}^{t}$ for all $i\in N$. Moreover, from the definition of $g(w^{0},s^{t},\mu^{t})]_{i}$ it is ensured that $s_{i}^{\max}$ is the fixed point of this gradient ascent update. Hence after a finite number of iterations, every agent reaches $s_{i}^{\max}$. Note that in this phase the model parameter $w^{0}$ remains unchanged.
In the second phase of the algorithm, we update the model parameter $w^{t}$. Note that since all the agents contribute the maximum amount of data they own, this phase is simply pure federated learning (without incentive design). As such, each agent now computes the local gradient $d_{i}^{t+1}$ and sends it to the center. The center then aggregates the gradients and takes a gradient ascent step with learning rate $\eta$. The center then broadcasts the updated parameter to the agents and the process continues.
We now provide the convergence guarantees of Algorithm˜3. Before that, let us discuss the necessary assumptions.
Assumption 4.
The smoothness and strong convexity assumptions have featured in several previous papers on FL (Karimireddy et al., 2020b; Yin et al., 2018). The main result of this section is as follows.
Theorem 3 (2P-UPBReD convergence to the optimal NE).
Suppose Assumptions  3 and 4 hold and we consider the utility function with payment scheme given in Equation˜15 with $\beta>\zeta+\tau$. Then, for every $w^{0}\in\mathbb{R}^{m}$ and $s^{0}\in\prod_{i\in N}S_{i}$, 2P-UPBReD (Algorithm˜3) converges to the Nash equilibrium $(w^{\texttt{{OPT}}},s^{\max})$ and is budget balanced. In particular, for any given $\epsilon>0$, we have
$\frac{1}{n}\sum_{i\in N}v_{i}(w^{\texttt{{OPT}}},s^{\max})-\frac{1}{n}\sum_{i\in N}v_{i}(w^{T},s^{\max})<\epsilon$
and $s^{T}=s^{\max}$ in $T=\kappa+T_{0}$ iterations provided we choose $\gamma=c$ (an universal constant) and $\eta=1/M$ with
$\kappa\geqslant\max_{i}\left\{\frac{\textstyle s^{\max}_{i}-s^{0}_{i}}{\textstyle c\,\Delta}\right\},$ (16)
$T_{0}>\left(\ln{\frac{f(w^{0},s^{\max})-f(w^{\texttt{{OPT}}},s^{\max})}{\epsilon}}\right)\Big/\left(\ln\frac{1}{{1-\frac{\nu}{M}}}\right)$ (17)
where, $f(w,s):=-\frac{1}{n}\sum_{i\in N}v_{i}(w,s)$ and $\Delta=\beta-\tau-\zeta$.
@@ 14      Send $d_{i}^{t+1}$ to the center
For the choice of $\beta>\zeta+\tau$ , we get $\Delta:=\beta-\zeta-\tau>0$. Applying the update rule for $s$ given by Algorithm˜3 in phase 1, we get
@@
Phase 2 For $t>\kappa$, we can focus on the second phase of the algorithm. From the fixed point property of $s^{\max}$, we can now focus entirely on the function $f(w,s^{\max})=-\frac{1}{n}\sum_{i\in N}v_{i}(w,s^{\max})$. From the algorithmic description, the second phase is just a simple gradient descent, run by the center, on $f(w,s^{\max})$. This is easy to see since at every iteration the center gets $d_{i}^{t+1}$ from all the agents $i\in N$, aggregates them and construct $\tilde{g}(w^{t},s^{\max})$, which is the gradient of $f(w,s^{\max})$ computed at $w^{t}$. We denote $f(w,s^{\max})$ with $f(w)$ in this part for notational cleanliness.
@@
which proves the theorem. Remark (Strong convexity and smoothness). The convergence rate is $\mathcal{O}(\ln{\nicefrac{{1}}{{\epsilon}}})$ which is the same as that of Algorithm˜1. However, Algorithm˜3 is cleaner in terms of the phases of convergence, the choices of the step sizes $\gamma,\eta$. The structural assumptions like strong convexity and smoothness may be relaxed at the expense of weaker rates. For example, we observe convergence rates of $\mathcal{O}(\frac{1}{\sqrt{\epsilon}})$ for non-convex smooth functions and $\mathcal{O}(\frac{1}{\epsilon})$ for convex smooth functions. ∎
Remark ($\gamma$ and rationality). Phase 1 of Algorithm˜3 achieves the optimal contribution $s^{\max}$ without updating the learning parameter $w$. The speed of convergence of this phase depends on the parameter $\gamma=c$, which can be interpreted as the minimum rationality level of the society. For instance, if all agents are sufficiently rational, i.e., $c$ is large, they may converge to $s^{\max}$ in a single iteration. However, the algorithm converges even if agents are boundedly rational. In Phase 2 of the algorithm, the agents stop updating their contributions and only update the gradients $d_{i}$ and the center accumulates them to update $w^{t}$.
Thanks to the smoothness and strong convexity properties of $f(\cdot,s^{\max})$, we can show that running Algorithm˜3 guarantees that the model iterate $w^{T}$ converges to $w^{\texttt{{OPT}}}$ at an exponential speed as well.
Corollary 1 (Iterate Convergence).
With the same setup as above and $\gamma=c$ (universal constant), $\eta=\frac{2}{M+\nu}$, we obtain $\|w^{T}-w^{\texttt{{OPT}}}\|_{2}<\epsilon,\text{ and }s^{T}=s^{\max}$, where $T=\kappa+\tilde{T}_{0}$, with the same $\kappa$ as in Equation˜16 and $\tilde{T}_{0}>\left(\ln{\frac{\|w^{0}-w^{\texttt{{OPT}}}\|_{2}}{\epsilon}}\right)/\left(\ln\frac{{1+\frac{\nu}{M}}}{{1-\frac{\nu}{M}}}\right)$.
@@
Remark (Truthfulness). An agent’s true type is $(s_{i}^{max},d_{i}^{t})$ in every iteration of phase $2$. We can use the payment scheme $p_{i}^{*}$ as defined in Equation˜13, with the necessary assumptions, in addition to the $z_{i}(s_{i},s_{-i})$ defined in this section. This modification will make the mechanism truthful with the trade-off that it will no longer be budget-balanced.
Remark (Robustness to adversarial behavior). In 2P-UPBReD, we have assumed that the agents report their gradients truthfully in phase $2$. If the agents misreport by sending arbitrary gradients not computed on any data, we have discussed one option of enforcing additional payment in the previous remark that compromises budget balance. If we do not wish to enforce additional payments, we can use some established methods in literature that are robust to adversarial agents (Yin et al., 2018; Karimireddy et al., 2020a; Ghosh et al., 2021). We describe one such method called trimmed mean  (Yin et al., 2018), in Section˜4.2 in the supplementary material and show how this method performs compared to UPBReD and 2P-UPBReD in the presence of adversarial agents in Section˜5.
4.2 Adversarial agents
@@
Let $\alpha$ be the fraction of adversarial agents in the system. Phase $1$ of the algorithm remains the same with agents first converging to $s^{\max}$. In the second phase which is a classic FedAvg algorithm the agents only report gradients computed on their training data. Instead of aggregating all gradients, for each coordinate of the gradients the values of the gradient corresponding to that coordinate are sorted. The lowest and highest $\alpha$ fraction of these values are discarded and the remaining gradient coordinates are used to update the parameter. We detail the algorithm in Algorithm˜4. In Section˜5 we show that this algorithm is robust to adversaries when the fraction of adversarial agents is small.
@@
Center broadcasts $w^{0},s^{0}$, set $t=0$
/* Begin phase 1 */
3 while $s\neq s^{\max}$ do
@@ 23 Center broadcasts $w^{t+1}=w$
5 Experiments
Our results in Sections˜3 and 4 guarantee convergence of Algorithms˜1 and 3 to a pure Nash equilibrium under the given assumptions. However, it is important to ask how these FL algorithms behave in practice with real datasets under more general setups. We capture results in these realistic settings in our experiments. We train models on the CIFAR-10  (Krizhevsky, 2009), FEMNIST, and Twitter datasets (Caldas et al., 2019) and compare the social welfare of UPBReD, 2P-UPBReD, and FedAvg. We also consider another algorithm FedAvgStrategic. In this algorithm, the agents choose to contribute their data strategically as in UPBReD, and stick to this contribution in the entire training phase of FedAvg (details in Algorithm˜5). The algorithm proposed in (Murhekar et al., 2023), which is closest to this paper, do not consider the simultaneous learning of parameters and strategy updates, and is therefore not meaningful to compare against.
@@ 16      Send $d_{i}^{t+1}$ to the center
The cost function used is linear $c_{i}(s_{i})=c_{i}\cdot s_{i}$, where $c_{i}$ is sampled from $U[0,1]$ for each agent. We choose our valuation function to be $v_{i}(w^{k},s^{k})=r_{i}-\mathcal{L}_{i}(w^{k},s^{k})$ where the $\mathcal{L}_{i}(w^{k},s^{k})$ is the loss evaluated on agents $i$’s test dataset at the $k$-th round of training. The social welfare is given by $\sum_{i\in N}v_{i}(w^{k},s^{k})$. We perform strategy updates using the numerical derivative as follows:
$s_{i}^{k+1}=s_{i}^{k}-\frac{L_{i}(w_{i}^{k})-L_{i}(w^{k-1})}{s_{i}^{k}-s_{i}^{k-1}}-c_{i}+\beta.$ (20)
The model parameters obtained by agent $i$ after performing one training step locally on the model initialized with $w^{k-1}$ is denoted by $w_{i}^{k}$. We use the cross-entropy loss function for $L_{i}$} on agent i’s dataset. All the experiments are performed on NVIDIA RTX A5000 and NVIDIA RTX A6000 with a 24GB and 48GB core respectively. The experiments required to obtain the results we report take several days to run. For comparing social welfare against $\beta$ and number of agents, we randomize over 10 runs, sampling an integer $s^{0}_{i}$ from $U[\nicefrac{{s^{\max}_{i}}}{{3}},\nicefrac{{2s^{\max}_{i}}}{{3}}]$ and cost from $U[0,1]$. We plot the average value of time taken across $10$ runs.
Dataset specifics. For CIFAR-10, we distribute the data equally among 100 agents by sampling the data uniformly at random. We train a CNN with 1,250,858 parameters for 100 rounds with agents updating their training datasets using the best response strategy in each round for UPBReD. We choose $\gamma=0.5$ and $\beta=2$. We use the Adam optimizer with learning rate $0.001$.
For FEMNIST, we use $20\%$ of the dataset that generates users with data that is not identically distributed as per the codebase provided by (Caldas et al., 2019). We train a CNN with $6,603,710$ parameters for $100$ rounds. We choose $\gamma=0.1$ and $\beta=2$. We use the Adam optimizer with learning rate $0.001$.
For the Twitter dataset, we use $15\%$ of the dataset that generates users with data that is identically distributed as per the codebase provided by (Caldas et al., 2019). We further group $1000$ agents together to create agents with larger data shares. We use an LSTM network to train for $50$ rounds. We choose $\gamma=0.5$ and $\beta=2$. We use the Adam optimizer with learning rate $0.00001$.
All results are summarized in Figure˜2.
(a) CIFAR-10
(b) FEMNIST
(c) Twitter
Figure 2: The three rows in this figure correspond to the experiments with the datasets CIFAR-10, FEMNIST, and Twitter respectively. The y-axis in every plot is the social welfare. The x-axis corresponds to the number of agents in the first plot, the choice of $\beta$ in the second, and the time taken (in seconds) in the third plot respectively in every row.
We observe that 2P-UPBReD and FedAvg have the same social welfare performance w.r.t. the number of agents (first column) and running times are nearly the same for all the datasets (third column). 2P-UPBReD outperforms UPBReD in nearly all the runs with respect to social welfare. We also observe that 2P-UPBReD converges to the maximum contribution of the agents for values of $\beta$ that are lower than the theoretically predicted ones. We also note that the data contributors get paid and data consumers pay in 2P-UPBReD and their payments are monotone non-decreasing with data contribution (Figure˜3). The utilities vary because the dataset on which the model is tested to obtain the accuracy varies for the agents.
In Section˜4.2, we describe the trimmed mean algorithm for 2P-UPBReD designed to be robust in the presence of adversarial agents. In this section, we demonstrate this robustness on the FEMNIST and CIFAR-10 dataset. We vary the adversarial fraction $\alpha\in\{0.025,0.05,0.075,0.1,0.125,0.15,0.175,0.2\}$, selecting suitable trimming parameters. Figure˜4 shows the ratio of social welfare with and without adversaries for UPBReD and 2P-UPBReD. While both suffer welfare loss under adversaries, applying the trimmed mean algorithm Algorithm˜4 keeps the welfare ratio close to $1$ for trimmed mean 2P-UPBReD.
(a) CIFAR-10
(b) FEMNIST
6 Conclusions and future work

### v4

@@ A.1 Proof of Theorem˜1
Using the update rules of Algorithm˜1 given by Equation˜3, we get
@@
A.2 Omitted examples and proofs of Section˜3.2
We perform some experiments on the CIFAR data set to show that agents have an incentive to misreport. Consider $2$ agents each having a set of $480$ data points. $u_{i}=56-L(w,s)-c_{i}(s_{i})$, where $L(w,s)$ is the cross entropy loss computed on the dataset. The costs for the agents are $c_{1}(s_{1})=0.2s_{i},c_{2}(s_{2})=0.1s_{2}$. When agents truthfully report $s_{i},d_{i}$ in UPBReD they observe a social welfare of $15.5$, with $s_{1}*=270,s_{2}*=165$ this yields a utility $u_{1}=510$. When agent misreports $s_{1}^{t}=240$ and reports $d_{1}^{t}=0_{m}$ in every training round yielding a utility of $u_{i}=64$. Clearly agent $1$ is better off by misreporting their type.
The algorithm to implement T-UPBReD is described below.
@@ 17     Send $\hat{v}_{i}^{t}=\hat{a}_{i}$
19 Center computes payments using Equation˜4 at $w^{t}$
@@
Proof of Theorem˜2
@@
Expanding the terms as defined in Equation˜4
@@
As social welfare is maximized by $a_{*}(\theta^{t})$
@@
Appendix B Appendix to Section˜4
B.1 Proof of Lemma˜1
@@
Using Equation˜5, we get the derivative of the utility function for $s_{i}\in(0,s_{i}^{\max})$ (note that this holds only for the interior of $S_{i}$, the derivative at the boundaries are zero by definition of function $g$ in Section˜3) to be $g(w,s,\mu)]_{i}=\frac{\partial}{\partial s_{i}}u_{i}(w,s)=\frac{\partial}{\partial s_{i}}v_{i}(w,s)-c_{i}^{\prime}(s_{i})+\beta$, where $c_{i}^{\prime}(s_{i})=\frac{d}{ds_{i}}c_{i}(s_{i})$. Using Assumption 3 and choosing in Equation˜5 the parameter $\beta>\zeta+\tau$, we have
@@
B.2 Proof of Theorem˜3
@@
For the choice of $\beta>\zeta+\tau$ , we get $\Delta:=\beta-\zeta-\tau>0$. Applying the update rule for $s$ given by Algorithm˜2 in phase 1, we get
@@
Phase 2: For $t>\kappa$, we can focus on the second phase of the algorithm. From the fixed point property of $s^{\max}$, we can now focus entirely on the function $f(w,s^{\max})=-\frac{1}{n}\sum_{i\in N}v_{i}(w,s^{\max})$. From the algorithmic description, the second phase is just a simple gradient descent, run by the center, on $f(w,s^{\max})$. This is easy to see since at every iteration the center gets $d_{i}^{t+1}$ from all the agents $i\in N$, aggregates them and construct $\tilde{g}(w^{t},s^{\max})$, which is the gradient of $f(w,s^{\max})$ computed at $w^{t}$. We denote $f(w,s^{\max})$ with $f(w)$ in this part for notational cleanliness.
@@
which proves the theorem. ∎
B.3 Proof of Corollary˜1
@@
B.4 Adversarial agents
@@
Let $\alpha$ be the fraction of adversarial agents in the system. Phase $1$ of the algorithm remains the same with agents first converging to $s^{\max}$. In the second phase which is a classic FedAvg algorithm the agents only report gradients computed on their training data. Instead of aggregating all gradients, for each coordinate of the gradients the values of the gradient corresponding to that coordinate are sorted. The lowest and highest $\alpha$ fraction of these values are discarded and the remaining gradient coordinates are used to update the parameter. We detail the algorithm in Algorithm˜4. In LABEL:app:experiments_adversarial we show that this algorithm is robust to adversaries when the fraction of adversarial agents is small.
@@
3 Center broadcasts $w^{0},s^{0}$, set $t=0$ while $s\neq s^{\max}$ do
@@ 23 Center broadcasts $w^{t+1}=w$
Appendix C Appendix to Section˜5
C.1 Algorithm for FedAvgStrategic
In the FedAvgStrategic algorithm, the agents choose to contribute their data strategically as in UPBReD, and stick to this contribution in the entire training phase of FedAvg. FedAvgStrategic is formally given by Algorithm˜5.
@@ 16      Send $d_{i}^{t+1}$ to the center
C.2 Model Architecture
We provide details of the model architecture used in our experiments in Tables˜2, 3 and 4.
Layer (type:depth-idx) Output Shape Param #
CNNModel [1, 10]
Conv2d: 1-1 [1, 32, 32, 32] 896
Conv2d: 1-2 [1, 32, 30, 30] 9,248
MaxPool2d: 1-3 [1, 32, 15, 15]
Conv2d: 1-4 [1, 64, 15, 15] 18,496
Conv2d: 1-5 [1, 64, 13, 13] 36,928
MaxPool2d: 1-6 [1, 64, 6, 6]
Linear: 1-7 [1, 512] 1,180,160
Linear: 1-8 [1, 10] 5,130
Total Params 1,250,858
Trainable Params 1,250,858
Non-trainable Params 0
Total Mult-Adds (MB) 20.83
Input Size (MB) 0.01
Forward/Backward Pass Size (MB) 0.70
Params Size (MB) 5.00
Estimated Total Size (MB) 5.71
Table 2: Layer-wise Summary of CNNModel for CIFAR
Layer (type:depth-idx) Output Shape Param #
CNNModel [1, 10]
Conv2d: 1-1 [1, 32, 28, 28] 832
MaxPool2d: 1-2 [1, 32, 14, 14]
Conv2d: 1-3 [1, 64, 14, 14] 51,264
MaxPool2d: 1-4 [1, 64, 7, 7]
Linear: 1-5 [1, 2048] 6,424,576
Linear: 1-6 [1, 10] 20,490
Total Params 6,497,162
Trainable Params 6,497,162
Non-trainable Params 0
Total Mult-Adds (MB) 17.15
Table 3: Layer-wise Summary of CNNModel used for FEMNIST
Layer (type:depth-idx) Output Shape Param #
Twitter [1, 2]
Embedding: 1-1 [1, 25, 300] 120,000,300
LSTM: 1-2 [1, 25, 100] 241,600
Linear: 1-3 [1, 128] 12,928
Linear: 1-4 [1, 2] 258
Total Parameters 120,255,086
Trainable Parameters 120,255,086
Non-trainable Params 0
Total Mult-Adds (MB) 126.05
Input Size (MB) 0.00
Forward/Backward Pass Size (MB) 0.08
Params Size (MB) 481.02
Estimated Total Size (MB) 481.10
Table 4: Layer-wise Summary of Twitter Model